### v1

# A Progressive Image Restoration Network for High-order Degradation Imaging in Remote Sensing

Abstract
Recently, deep learning methods have gained remarkable achievements in the field of image restoration for remote sensing (RS). However, most existing RS image restoration methods focus mainly on conventional first-order degradation models, which may not effectively capture the imaging mechanisms of remote sensing images. Furthermore, many RS image restoration approaches that use deep learning are often criticized for their lacks of architecture transparency and model interpretability. To address these problems, we propose a novel progressive restoration network for high-order degradation imaging (HDI-PRNet), to progressively restore different image degradation. HDI-PRNet is developed based on the theoretical framework of degradation imaging, offering the benefit of mathematical interpretability within the unfolding network. The framework is composed of three main components: a module for image denoising that relies on proximal mapping prior learning, a module for image deblurring that integrates Neumann series expansion with dual-domain degradation learning, and a module for super-resolution. Extensive experiments demonstrate that our method achieves superior performance on both synthetic and real remote sensing images.
Index Terms:
High-order degradation model, image restoration, remote sensing image, deep unfolding network.
I Introduction
Remote sensing images contained abundant spatial texture details and scene semantic information, which have been widely applied in tasks such as scene segmentation [1]. However, due to the complex environment of remote sensing imaging and the inherent limitations of sensor hardware, the observed remote sensing images inevitably degrade. The low quality of remote sensing images limits their practical application in high-level remote sensing tasks directly. To address this problem efficiently with minimal cost, employing image restoration techniques is optimal, as it avoids the need for upgrading equipment. These techniques reconstruct a high-resolution (HR) image, enriched with detailed high-frequency textures, from a low-resolution (LR) image. As a result, image restoration has become increasingly popular in the fields of remote sensing and computer vision.
The classical first-order image degradation model can be expressed mathematically as follows
$\begin{split}\bm{g}&=(\bm{h}*\bm{u})_{\downarrow_{\bm{s}}}+\bm{n},\end{split}$ (1)
where $\bm{g}$ is the observational LR image, $\bm{u}$ is the HR image with sharp details, $\bm{h}$ is the blur kernel, $\bm{n}$ is the additive white gaussian noise (AWGN), and * denotes the 2-D convolution operator, $\downarrow_{\bm{s}}$ denotes the downsampling operator with scale factor $\bm{s}$.
The restoration of degradation images in remote sensing is a challenging task, as it involves restoring the sharp HR image $\bm{u}$ from the deteriorated LR image $\bm{g}$. This problem (1) is considered ill-posed because there are multiple HR solutions that can map to the same LR image, resulting in a non-unique solution. Various methods [2, 3, 4] for image restoration have been investigated. Although many algorithms ignore blur and noise degradation in the forward model (1) and rely solely on the ideal bicubic operator to simulate the super-resolution degradation model, these approaches do not align with the actual physical imaging mechanism of remote sensing, rendering these algorithms impractical in real-world remote sensing applications [5, 6].
Recently, numerous deep learning-based methods have been proposed for image restoration [7, 8, 9, 10, 11, 12, 13, 14] and have demonstrated strong performance across various datasets and evaluation metrics. These methods aimed to integrate advanced deep learning modules and techniques such as channel attention mechanisms [10], spatial attention mechanisms [11], transformer mechanisms [12], and generative adversarial networks (GANs) [15]. The primary aim of these methods is to learn the mapping between LR and HR images. However, several challenges remain. 1) Most super-resolution algorithms use the first-order degradation model, which is inadequate for real images as these are often affected by high-order complex degradations; 2) the end-to-end black-box nature of deep learning-based algorithms makes it challenging to distinguish how the network architecture influences its prediction, emphasizing the need for enhanced transparency and interpretability in these frameworks; 3) many deep learning techniques focus on designing sophisticated network modules and depend on expanding them deeper and wider to enhance performance, this results in increased computational needs and challenges related to the smart design of network architectures.
To address these problems, we introduce a high-order degradation model to describe the combined effects of various complex degradations in remote sensing imaging, including atmospheric transmission degradation, sensor imaging system degradation, satellite platform flutter degradation, etc. Based on the high-order degradation imaging, we propose a new high-order progressive restoration network, called HDI-PRNet, which aims to recover multiple layers of degradation. Specifically, each stage of this model is composed of three components: a module for image denoising that relies on proximal mapping prior learning, a module for image deblurring that integrates Neumann series expansion with dual-domain degradation learning, and a module for super-resolution. The main contributions in this paper can be summarized as follows.
  • We introduce a novel high-order progressive remote sensing restoration network, termed HDI-PRNet, constructed around a high-order backward degradation model. Our framework adeptly manages complex degradation mechanisms, notably enhancing restoration performance, particularly in real remote sensing applications. The structures aligning backward degradation tasks with respective modules render HDI-PRNet more transparent and interpretable.
  • A denoising module is developed using proximal mapping operator learning. By leveraging the multi-scale encoder-decoder learning and attention mechanisms, it emphasizes crucial denoising prior details required at each stage while preserving fine detail information.
  • We design a deblurring module that leverages the Neumann series expansion in combination with a dual-domain backward degradation learning block. This approach efficiently acquires prior knowledge of blur degradation operators in both spatial and frequency domains, leading to the integration of dual-domain representations.
  • Experimental results confirm that the HDI-PRNet framework can obtain high-quality restored images on various randomly degraded synthetic low-resolution images and real remote sensing images.
This paper is organized as follows. Section II covers related work in image restoration. Our proposed HDI-PRNet framework is detailed in Section III. we analyze our experimental results and ablation experiments in Section IV. Finally, Section V presents the conclusions.
II Related Work
In this section, we briefly review the related works on natural image restoration algorithms, remote sensing image restoration algorithms, and multiple degradation models.
II-A Image restoration for natural images
Recently, with the rapid development of neural network technology, deep learning methods have achieved excellent performance in the field of computer vision. Dong et al. [16] proposed SRCNN, which first applied a three-layer convolutional neural network for image super-resolution (SR). Zhang et al. [17, 18] proposed the residual dense block (RDB), which fully utilizes all of the layers in it through local dense connections, and proposed the residual dense network (RDN). Zhang et al. [10] proposed the structure of residuals in residuals (RIR) to form very deep networks, and also proposed the residual channel attention network (RCAN) for image SR. Zamir et al. [19] introduced Restormer, a transformer model that performs well in a variety of image restoration tasks. Meanwhile, Fan et al. [4] developed an incremental learning framework, called Nest-DGIL, using Nesterov’s method to reconstruct images from compressed sensing data. Dai et al. [20] proposed a high-order attention network for single image SR (SAN). Liang et al. [12] proposed an image restoration model based on the Swin Transformer (SwinIR), and achieved good performance in image SR task, image denoising and JPEG artifact removal tasks. Mei et al. [21] proposed a novel non-local sparse attention (NLSA) with dynamic sparse attention pattern for image SR. Cui et al. [8] introduced FSNet, an image restoration network focusing on frequency information selection. Chen et al. [9] introduced a nonlinear activation-free network (NAFNet), which offers a simplified and efficient baseline for recovery models by replacing the activation function with multiplication. Chen et al. [22] proposed the dual aggregation transformer (DAT), which aggregates spatial-channel features in the interblock and intra-block, achieving powerful feature representation capabilities and achieving fascinating performance on SR task. While these algorithms have demonstrated fascinating results in natural image restoration tasks, they often prioritize the development of complex high-performance learning modules and neglecting the interpretability of the network structure.
II-B Image restoration for remote sensing images
Nowadays, with the increasing application of remote sensing images in various fields, the remote sensing image restoration algorithms have attracted increasing attention. They can improve the quality of remote sensing images, restoring more distinct details compared to the original low-quality images, without equipment upgrades.
With the development of deep learning and the success of natural image processing algorithms, lots of remote sensing image restoration algorithms have been proposed. Lei et al. [23] designed a combined local-global network (LGCNet), which learns multilevel feature representations of images through different convolutional layers and global-local features of ground objects and environmental priors, to reconstruct high-quality remote sensing images. Zhang et al. [24] proposed a kind of novel mixed high-order attention network (MHAN) to fully exploit hierarchical features for detail restoration. Xu et al. [25] designed a global-local fusion-based cloud removal algorithm (GLF-CR) that combines synthetic aperture radar and optical image feature information for remote sensing image cloud removal. Xiao et al. [26] designed a top-k token selective transformer (TTST) for remote sensing image super-resolution to flexibly filter information in self-attention. Wang et al. [27] proposed a novel U-shaped network based on hybrid attention (HAUNet), which is designed based on a multiscale structure and a hybrid convolutional attention mechanism to enhance the representation of global features. Lei et al. [28] proposed the transformer-based enhancement network (TransENet), which utilizes the transformer block to extract the high/low-dimensional features to enhance the representation of high dimensions features. TransENet performs well on the synthetic UCMecred and AID dataset.
However, most of these image restoration algorithms use simple bicubic downsampling to simulate remote sensing image degradation, which is inconsistent with real remote sensing degradation. This is one of the reasons why the algorithms perform poorly on real remote sensing images. In order to solve this problem, we design a network based on a high-order degradation model that is closer to real degradation.
II-C Multiple degradation models
The first-order degradation model (1) is widely used in image restoration algorithms and has achieved good performance. Most of these algorithms simplify degradation into a simple bicubic downsampling process. However, real-world imaging degradation is a complex, multi-faceted process, and simple bicubic downsampling limits the robustness of the algorithm and does not accurately simulate real-world image degradation. Wang et al. [29] proposed a high-order degradation model that simulates complex real-world degradation by combining multiple degradations (including blur, noise, downsampling, etc.), and achieved excellent real-world super-resolution results. Zhang et al. [30] proposed a complex degradation model which consists of multiple random shuffled degradations. Potlapalli et al. [31] designed a prompt learning restoration network (PromptIR) for different types of degradation, providing a versatile and efficient plug-in module with few lightweight prompts that can be used to restore images of various types and degradation levels. Kong et al. [32] proposed a sequential learning strategy and a prompt learning strategy for all-in-one image restoration. These two strategies have good effects on the CNN and Transformer backbones, and they can promote each other to learn effective image representations. In comparison, our work is focused on developing an unfolding network for real remote sensing image restoration that is conceptually simple but more interpretable.
Figure 1: The overall framework of the proposed HDI-PRNet solving the high-order degradation model.
III Methodology
This section first briefly introduces the high-order degradation model, followed by a detailed explanation of the proposed progressive restoration architecture designed to solve high-order degradation imaging. We also introduce a multi-supervising loss function to train the proposed architecture.
III-A High-order Degradation Model
An image degradation model is formulated to simulate the mechanisms that lead to a decrease in image quality [33, 34]. The classical first-order image degradation (1) is rewritten as
$\bm{g}=\mathcal{D}(\bm{u})=(\bm{h}*\bm{u})_{\downarrow_{\bm{s}}}+\bm{n},$
where $\mathcal{D}$ denotes the degradation operator, it includes factors such as the blur, downsampling, and noise degradation. However, the first-order model described above remains inadequate for addressing certain complicated real-world degradation, particularly unknown noises and complex artifacts. The authors in [29] improve the first-order degradation model by extending it to a more flexible higher-order degradation to better represent real-world degradations, as defined by
$\bm{g}=\mathcal{D}_{k}(\cdots\mathcal{D}_{2}(\mathcal{D}_{1}(\bm{u}))),$ (2)
where $\mathcal{D}_{\ell}$ denotes the $\ell$-th imaging sub-system or degradation sensor ($1\leq\ell\leq k$).
In this work, we decompose the $k$-order degradation framework into a series of $k$ sequential first-order models, described as follows
$\begin{split}\bm{g}_{1}=&\mathcal{D}_{1}(\bm{u})=\mathcal{N}_{1}(\mathcal{S}_{1}(\mathcal{B}_{1}(\bm{u})))\\ \bm{g}_{2}=&\mathcal{D}_{2}(\bm{g}_{1})=\mathcal{N}_{2}(\mathcal{S}_{2}(\mathcal{B}_{2}(\bm{g}_{1})))\\ &\cdots\\ \bm{g}:=&\bm{g}_{k}=\mathcal{D}_{k}(\bm{g}_{k-1})=\mathcal{N}_{k}(\mathcal{S}_{k}(\mathcal{B}_{k}(\bm{g}_{k-1}))),\end{split}$ (3)
where the functions $\mathcal{B}_{k}(\cdot)$, $\mathcal{S}_{k}(\cdot)$, $\mathcal{N}_{k}(\cdot)$ represent the blur, downsampling, and noise degradation operators, respectively. The formula (3) outlines a hierarchical degradation model that provides a more precise representation of the real remote sensing imaging system.
III-B Algorithm
The restoration algorithm based on Eq. (3) can be divided into $k$ stages to be solved progressively, where $k$ stages correspond to the high-order degradation mechanism. We assume $\bm{g}_{0}=\bm{u}$, $\bm{g}_{\ell}=\mathcal{D}_{\ell}(\bm{g}_{\ell-1})$ ($1\leq\ell\leq k$) and $\bm{g}_{k}=\bm{g}$. The energy minimization model at $\ell$-stage can be written as
$\bm{g}_{\ell-1}=\underset{\bm{g}}{\arg\min}\|\bm{g}_{\ell}-\mathcal{S}_{\ell}(\mathcal{B}_{\ell}(\bm{g}))\|_{2}^{2}+\lambda_{\ell}f_{\ell}(\bm{g}),$ (4)
where the regularization $f_{\ell}(\cdot)$ can be used to constrain the geometric prior of a given image in denoising subproblem.
There have been several excellent explorations in solving (4). One of main categories involves step-by-step degradation representations and typically consists of three components: prior denoising or local filtering, super-resolution or interpolation, and blind degradation prediction. The above three components are performed either separately (iteratively) or jointly. Therefore, the above energy minimization model (4) can be divided into three sub-problems, as follows
$\bm{g}^{dn}_{\ell}$ $=\underset{\bm{g}}{\arg\min}\|\bm{g}-\bm{g}_{\ell}\|_{2}^{2}+\lambda_{\ell}f_{\ell}(\bm{g}),$ (5)
$\bm{g}^{sr}_{\ell}$$\in\{\bm{g}|\mathcal{S}_{\ell}(\bm{g})=\bm{g}^{dn}_{\ell}\},$ (6)
$\bm{g}_{\ell-1}$ $:=\bm{g}^{db}_{\ell}\in\{\bm{g}|\mathcal{B}_{\ell}(\bm{g})=\bm{g}^{sr}_{\ell}\}.$ (7)
The classical solution methods of the above subproblems depend on the types and levels of degradation, and the forward simulation typically involves straightforward synthetic degradations. Furthermore, any imprecise estimations of degradation in real images will unavoidably lead to artifacts. Therefore, the following literature will focus on developing a progressive unfolding restoration network for real remote sensing imaging.
III-C Proposed Framework
III-C1 Overall Architecture
Deep neural networks with their powerful non-linear representation capabilities have become increasingly popular in addressing reconstruction challenges associated with degraded data. Building upon the progressive operator splitting theory described above, we construct a deep unfolding network for the degradation reconstruction. The overall framework we propose is shown in Fig. 1(a).
The input variable $\bm{g}_{\ell}$ at $\ell$-stage first undergoes a deep prior unfolding denoising sub-network to obtain the clean approximation $\bm{g}^{dn}_{\ell}$. In addition, the super-resolution distillation SR module is proposed to exploit the structural characteristics among different features by interpolating multiscale representations, thereby improving the resolution of an image $\bm{g}^{dn}_{\ell}$ and improving image quality. This super-resolution estimation $\bm{g}^{sr}_{\ell}$ is then refined through a deep blind-deblurring subnetwork, where each block corresponds to one Neumann expansion term of the blind-deblurring sub-problem. The initial reconstruction $\bm{g}^{sr}_{\ell}$ is processed through multiple blocks of the sub-network to achieve the final deblurring reconstruction. Specifically, our innovative network design is carried out hierarchically. The framework consists of $k$ stages corresponding to the $k$-order degradation model, and each stage consists of three modules corresponding to three sub-problems (5)-(7).
To better leverage the prior information, we employ one proximal mapping subnetwork to independently solve (5), represented as the denoising module. For the subproblem of blind deblurring, we improve the feature representation capability of the network by learning truncated Neumann series to replace the direct inverse operator $\mathcal{B}^{-1}$. This structure ensures the robustness of the framework in handling multiple or different types of degraded image. Each degradation sub-system will solve the inverse problem progressively. Additionally, to reduce information loss between network stages, we add intermediate losses to ensure that the submodule learns the corresponding subproblem, to improve the spatio-temporal feature representation of the framework, and also to ensure the interpretability of the algorithm. The specific design of each sub-network will be introduced in the following subsections.
III-C2 Image denoising proximal mapping module
The sub-problem (5) is a classical denoising optimization, which is highly effective in removing various types of noise. The solution of the optimization (5) with the general regularization term $f_{\ell}$ can be given by the proximal operator as
$\bm{g}^{dn}_{\ell}$ $=\textbf{prox}_{\lambda_{\ell},f_{\ell}}(\bm{g}_{\ell}).$ (8)
Since $f_{\ell}(\cdot)$ is the geometric prior term of the solution image $\bm{g}$, hence one of the key advantages of the sub-problem (5) is its ability to preserve important image details. However, the regularization parameter $\lambda_{\ell}$ in (5) controls the trade-off between the fidelity to the observed data and the smoothness of the image. If $\lambda_{\ell}$ is larger, the denoised image is imposed stronger smoothing, leading to a potentially smoother but a loss of fine-grained details like edges, textures, and small structures. If $\lambda_{\ell}$ is smaller, then the denoising effect may be weaker, resulting in an image closer to the original but potentially with more noise.
The proximal mapping module can analyze the local neighborhood of each pixel and fuse features to reduce the noise component while preserving the details of the image [35]. The optimization problem (5) provides a theoretical guarantee for the performance and convergence of the proximal mapping module. In other words, it typically operates by solving a proximal operator problem that seeks to minimize the denoising problem (5). Compared to some complex iterative denoising methods that require a large number of calculations for each pixel and multiple iterations over the entire image, the proximal mapping module can achieve good denoising results with fewer computational resources [36, 4]. This makes it suitable for real-time or near-real-time applications such as video denoising or high-throughput image processing in industrial inspection.
To obtain a more accurate prior denoising expression of (8), we design an image denoising unfolding module to learn the proximal mapping, as shown in Fig. 1(b). The proximal mapping $\textbf{prox}_{\lambda_{\ell},f_{\ell}}$ is designed as a coarse to fine architecture, uses an encoder and a decoder to learn the generalized multiscale prior $\lambda_{\ell}f_{\ell}(\bm{g})$, and then obtains the solution of (8). Firstly, the denoising module adopts a convolutional design to extract shallow prior features. Secondly, it involves three scales, and we use Residual Channel Attention Blocks (RCAB) to extract and fuse multi-scale prior features at three scales (see Fig.1(b) and (g)). The skip connections between downsampling and upsampling at each scale are also processed using RCAB, and the channel number of the three scales is set to 120, 140, and 160 at resolutions $1,1/2,1/4$, respectively. Then, we introduce a skip connection before and after the encoder decoder to ensure the stability of proximal mapping learning. Finally, in terms of sampling operators, we use bilinear sampling to perform down-sampling and up-sampling operations.
III-C3 Image super-resolution module
Image super-resolution is a method employed to improve the resolution of an image, resulting in a clearer and more detailed appearance. This enhancement is typically accomplished using algorithms and models that estimate high-resolution predictions from low-resolution inputs [37].
The CNN image super-resolution leverages deep learning techniques to enhance the resolution of images. It typically involves training a CNN model on low-resolution and high-resolution image pairs, where the network learns to reconstruct high-resolution images from their low-resolution counterparts. The proposed framework employs an SR module with convolution and pixelshuffle to solve the problem at $\ell$-stage as
[…]
Since the proposed framework is a $k$-stage cascaded architecture that processes sub-problems progressively, the first-stage SR module resizes the image from the LR space to the middle resolution space, and then the final stage resizes it to the HR space. In this case, we need to perform SR tasks with a different upsampling ratio ($\uparrow_{\bm{s}}$) in the $k$ stages. To achieve efficiently fractional super-resolution without introducing new degradation, we simply use convolution and bilinear interpolation to perform upsampling operations to solve (6), as shown in Fig. 1(c). This module can ensure that the image will not have problems such as checkerboard artifacts caused by transposed convolution. However, such an SR module can efficiently adjust the spatial resolution. Although bilinear interpolation is considered to have defects in sharpening edge details, the deblurring problem can be handled by the subsequent deblurring module.
III-C4 Image deblurring module
Let the blurring operator $\mathcal{B}_{\ell}$ be a bounded linear operator that satisfies $\rho(\mathcal{I}-\mathcal{B}_{\ell})<1$ on a Banach space $\mathcal{X}$. The Neumann expansion series of $\mathcal{B}_{\ell}^{-1}$ is given by the following infinite series
$\begin{split}\mathcal{B}_{\ell}^{-1}=\left(\mathcal{I}-(\mathcal{I}-\mathcal{B}_{\ell})\right)^{-1}=\sum_{i=0}^{+\infty}(\mathcal{I}-\mathcal{B}_{\ell})^{i},\end{split}$ (9)
where $\mathcal{I}$ is an identity operator, $(\mathcal{I}-\mathcal{B}_{\ell})^{0}=\mathcal{I}$ (the identity operator) and $(\mathcal{I}-\mathcal{B}_{\ell})^{i}=(\mathcal{I}-\mathcal{B}_{\ell})\circ\cdots\circ(\mathcal{I}-\mathcal{B}_{\ell})$ ($i$ times composition of $(\mathcal{I}-\mathcal{B}_{\ell})$ with itself for $i\geq 1$). To facilitate calculations, we use a $m$-term truncated Neumann series expansion to approximate $\mathcal{B}_{\ell}^{-1}$ as follows
$\begin{split}\bm{g}_{\ell}^{db}=\mathcal{B}_{\ell}^{-1}(\bm{g}^{sr}_{\ell})&\approx\sum_{i=0}^{m}(\mathcal{I}-\mathcal{B}_{\ell})^{i}(\bm{g}^{sr}_{\ell}).\end{split}$ (10)
Even if the condition $\rho(\mathcal{I}-\mathcal{B}_{\ell})<1$ is not met, the truncated Neumann series (10) can still approximate the inverse operator $\mathcal{B}_{\ell}^{-1}$, providing a form of implicit regularization [38].
To obtain accurately blur degradation, we use the Dual Domain Degradation Learning Block (DDLB) designed based on the ResFFT-ReLU Block [39] as follows
$\begin{split}\mathcal{I}-\mathcal{B}_{\ell}=\mathcal{I}+\textbf{DEA}(\mathcal{B}^{Spatial}_{\ell},\mathcal{B}^{Frequency}_{\ell})\end{split}$ (11)
to learn (10), as shown in Fig. 1(d) and (h). Here, we split the blur operator $\mathcal{B}_{\ell}$ into a multi-scale dual-domain feature representation. Therefore, DDLB includes the spatial transformation branch $\mathcal{B}^{Spatial}_{\ell}$, the frequency transformation branch $\mathcal{B}^{Frequency}_{\ell}$, the dual-domain expression aggregation (DEA) and the identity transformation $\mathcal{I}$.
In the spatial domain branch, the Conv-PRelu-Conv configuration is used to learn and simulate the degradation process $\mathcal{B}^{Spatial}_{\ell}$. This structure is widely adopted in CNN architectures and has been validated to effectively capture local spatial features. Regarding the branch in the frequency domain, the image is transformed into a complex tensor $\bm{z}=\bm{m}e^{j\bm{\beta}}$ by a Fourier transform, which is divided into two components: amplitude $\bm{m}$ and phase $e^{j\bm{\beta}}$. The amplitude encodes the image’s color information, whereas the phase represents the texture’s positional information. Similarly to the spatial domain strategy, the Conv-PRelu-Conv block is used to model and simulate the degradation mechanism $\mathcal{B}^{Frequency}_{\ell}$. The PRelu function helps in phase selection as it satisfies $\textbf{PRelu}(\bm{m}e^{j\bm{\beta}})=\bm{m}\textbf{PRelu}(e^{j\bm{\beta}})$. The first convolution step provides greater flexibility in phase selection, while the last convolution consolidates the selected phases to accurately express the blur operator $\mathcal{B}^{Frequency}_{\ell}$.
The representations of degradation, $\mathcal{B}^{Spatial}_{\ell}$ in the spatial domain and $\mathcal{B}^{Frequency}_{\ell}$ in the frequency domain, are not adequately integrated when combined solely with an addition operator. To enhance the interaction and blending of spatial and frequency domain representations, we developed the dual-domain expression aggregation (DEA), which incorporates a channel-interaction aggregation (CIA) module and a spatial-interaction aggregation (SIA) module, positioned between the spatial and frequency domain branches, illustrated in Fig. 1 (e)-(f). The application of CIA and SIA enables the adaptive synchronization of dual domain degraded features, which mutually improve their insights, refine the representation of the features and promote the integration of the degraded information.
Figure 2: Visualization comparison for different intermediate loss weight $\alpha$.
III-D Loss Function
The HDI-PRNet is composed of $k$ stages, with each stage containing three sub-modules. To optimize the parameters of these modules, we minimize the total loss, represented as
$\begin{split}\mathcal{L}_{Total}=\mathcal{L}_{Rec}+\alpha\mathcal{L}_{Int},\end{split}$ (12)
which consists of two components: the image reconstruction loss $\mathcal{L}_{Rec}$ and the intermediate loss $\mathcal{L}_{Int}$. The weight parameter $\alpha$ is set to 1.0.
III-D1 Image reconstruction loss
To achieve high-quality reconstruction of the degraded image, the $\ell_{2}$ loss function is employed to measure the difference between the restored image $\bm{g}_{0}$ and the HR ground-truth $\bm{u}_{HR}^{gt}$. It is represented as:
$\begin{split}\mathcal{L}_{Rec}=\|\bm{g}_{0}-\bm{u}_{HR}^{gt}\|_{2}.\end{split}$ (13)
III-D2 Intermediate loss
In deep-supervising learning, intermediate loss assists in training by offering additional guidance in various stage modules. This loss can improve feature representation, accelerate convergence, and improve the model’s predictive abilities. To verify that each module is executing its designated function, we incorporated an intermediate loss defined as follows:
$\begin{split}\mathcal{L}_{Int}&=\sum_{i\in\mathcal{A}}\|\bm{g}_{i}-\bm{u}_{i}^{gt}\|_{2},\end{split}$ (14)
where $\mathcal{A}$ represents the subscripts for the intermediate predictions of HDI-PRNet. In particular, the intermediate loss regulates distinct modules responsible for specific learning activities, each with a clear objective, which significantly enhances the interpretability of the proposed network.
IV Numerical Experiments
IV-A Implementation Details
IV-A1 Datasets
In our experimental investigation of different restoration methods for remote sensing, we have chosen to focus on well-known datasets, namely AID [40], WHU-RS19 [41], and WHU Building [42].
  • 1)
    AID dataset [40] consists of 10,000 images, each measuring $600\times 600$ pixels with a spatial resolution of 0.5 m/pixel (0.5 meters per pixel). This collection represents 30 different land-use categories, such as airports, farmlands, beaches, deserts, etc.
  • 2)
    WHU-RS19 dataset [41] consists of 1005 remote sensing images, each $600\times 600$ pixels in size, with a maximum spatial resolution of 0.5 m/pixel. The dataset represents 19 unique land types, such as airports, bridges, mountains, and rivers, among others.
  • 3)
    WHU Building dataset [42] features more than 220,000 individual buildings derived from aerial imagery at a spatial resolution of 0.075 m in Christchurch, New Zealand. The WHU Building dataset contains 4736 and 1036 aerial images with $512\times 512$ resolution for the training and validation sets.
For training and assessment of our methodology, the AID and WHU Building sets are used for training, while WHU-RS19 serves as a test set.
TABLE I: Ablation study of model degradation orders.
Network type 1st-order degradation images 2nd-order degradation images 3rd-order degradation images Params (M) FLOPs (G)
PSNR (dB) SSIM PSNR (dB) SSIM PSNR (dB) SSIM
1st-order network 31.80 0.8492 28.81 0.7438 27.20 0.6776 8.19 167.37
2nd-order network 31.70 0.8467 30.03 0.7931 28.49 0.7350 16.39 338.72
3rd-order network 31.48 0.8428 29.95 0.7913 28.64 0.7408 24.58 515.43
IV-A2 High-order Degradation Simulation
Inspired by the image degradation mechanism as detailed in [29], we employ a high-order degradation model to simulate real degradation on hyper-resolution datasets. Initially, the blurring model is used on the original image, followed by applying the undersampling operator to the blurred image. Lastly, the noise model is applied to the image that has been both undersampled and blurred. We continuously implement these degradation models to simulate the $k$-order hybrid image degradation (see Fig. 1(a) for more details).
To simulate the $\ell$-th blur degradation $\mathcal{B}_{\ell}(\cdot)$, we configure the blur kernel size as $12n+$ 1 where $n$ is randomly selected from the set $\{3,4,\dots,10\}$. We employ the selection strategy to determine the blur kernel type: 1) the sinc filter is chosen with a probability of $0.1$ to simulate the ringing and overshoot effects in images; 2) the isotropic Gaussian, anisotropic Gaussian, isotropic generalized Gaussian, anisotropic generalized Gaussian and isotropic / anisotropic plateau-shaped blur kernels are generated with a probability of $0.9$, following proportions of 0.45:0.25:0.12:0.03:0.12:0.03. The Gaussian blur standard deviation is selected between $0.1$ and $2.0$, but for the second degradation stage, this range is narrowed to $0.1$ to $1.0$. The shape parameters for the generalized Gaussian and plateau-shaped kernels are chosen from $0.5$ to $4$ and $1$ to $2$, respectively. The second stage of blur degradation is omitted with a probability of $0.2$.
To simulate the $\ell$-th resizing degradation $\mathcal{S}_{\ell}(\cdot)$, we perform upsampling, downsampling, or maintain the original image resolution with probabilities of 0.2, 0.7, and 0.1, respectively (reconfigured to probabilities of 0.3, 0.4, and 0.3 during the second degradation stage). The resizing scale is randomly chosen between 0.5 and 1.5 (ranging between 0.8 and 1.2 for the second degradation step). The sampling method is randomly selected from bilinear interpolation, bicubic interpolation, and area resizing. Finally, we resize the images to achieve the desired low-resolution.
To simulate the degradation of noise pollution $\mathcal{N}_{\ell}(\cdot)$ at $\ell$-th stage, Gaussian and Poisson noise are incorporated, each with a probability of $0.5$. The level of Gaussian noise ranged from $1$ to $25$, and the level of Poisson noise from $0.05$ to $2.5$. In contrast, the noise level used in the second degradation stage ranges from $1$ to $20$ for Gaussian and $0.05$ to $2.0$ for Poisson. Furthermore, gray noise was applied with a probability of 0.4.
IV-A3 Evaluation Metrics
We evaluate the image restoration in remote sensing using peak signal-to-noise ratio (PSNR), structural similarity (SSIM) [43], and the blind/referenceless image spatial quality evaluator (BRISQUE) [44]. The superior restoration performance is reflected in elevated PSNR and SSIM values, along with reduced BRISQUE values.
Figure 3: Visualization comparison for super-resolution scales x2, x3 and x4 on WHU-RS19 dataset.
IV-A4 Compared Methods
We conduct a group of studies with fourteen other methods to evaluate the performance of the proposed method. We compare our MVMS-RCN with general image restoration approaches such as RCAN [10], SwinIR [12], RDN [17], SAN [20], DAT [22], D-DBPN [45], CAT [46], and RGT [47], as well as super-resolution methods in remote sensing, such as MHAN [24], TTST [26], HAUNet [27], TransENet [28], CTNet [48], and HSENet [49].
IV-A5 Training Settings
We use a batch size of $32$ to train our model. Each input of the LR image is randomly cropped to dimensions of $64\times 64$, with a total of $200K$ training iterations. The training patches undergo augmentation via random horizontal flips and rotations. To keep fair comparisons, we use the Adam optimizer with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.99$. The initial learning rate is set as $2\times 10^{-4}$, which is halved at all milestones of $[20K,120K,160K,180K]$. Our models are implemented using PyTorch and run on a Nvidia A100 GPU.
TABLE II: Comparison for different loss weight parameters $\alpha$ on AIRS dataset.
Loss weight $\alpha$ PSNR (dB) SSIM
0 29.96 0.7915
0.1 30.06 0.7933
0.5 30.04 0.7934
1.0 30.03 0.7931
TABLE III: Ablation study of the denoising module configuration.
Scale number Block number Params (M) FLOPs (G) PSNR (dB) SSIM
1 2 3 1 2
7.75 279.94 29.93 0.7893
10.04 295.19 29.97 0.7904
11.98 298.64 29.98 0.7923
16.39 338.72 30.03 0.7931
TABLE IV: Ablation study for different orders of the Neumann series expansion in deblurring module.
Orders number Params (M) FLOPs (G) PSNR (dB) SSIM
1 3 5
13.37 208.55 29.83 0.7902
14.88 273.64 29.95 0.7926
16.39 338.72 30.03 0.7931
TABLE V: Ablation study of dual-domain degradation learning block.
SB FB DEA Params (M) FLOPs (G) PSNR (dB) SSIM
15.21 308.72 29.83 0.7884
13.77 205.76 29.90 0.7892
16.36 338.47 30.00 0.7933
16.39 338.72 30.03 0.7931
IV-B Ablation Study
We conduct extensive ablation studies to evaluate the contributions of each component within the proposed HDI-PRNet framework. These experiments involve analyzing the advantages of high-order networks, the loss weight, the denoise module architecture, the number of Neumann series expansion orders, and the effectiveness of DDLB.
IV-B1 Evaluation for high-order degradation
To demonstrate the advantages of high-order framework, we perform ablation studies, as shown in Table I. Our analysis assesses the performance of framework with differing degradation orders across various degraded images to demonstrate their robustness. We can see from Table I that the low-order degradation network exhibits significantly decreasing performance when faced with complex high-order degraded images. In contrast, high-order degradation network achieves more consistent and superior results in different degradation images, underscoring their increased robustness. For optimal configuration, we chose the second-order degradation architecture because it provides a suitable balance between network complexity and reconstruction performance.
IV-B2 Sensitivity of the loss weight $\alpha$
The WHU-RS19 dataset is used to explore the optimal value for the loss weight $\alpha$. As shown in Table II, setting the weight $\alpha=0$ results in inferior image reconstruction performance compared to the methods where $\alpha\neq 0$. This indicates that the intermediate loss that supervises each submodule is advantageous.
Fig. 2 illustrates the intermediate predictions and the final restored results with different loss weights. Evidently, when the weight $\alpha$ is set to 00, each sub-module predicts an approximate black-box image, obscuring the understanding of the individual roles and impacts of each sub-module. On the other hand, incorporating the intermediate loss illustrates that the sub-modules are beneficial to their designated sub-tasks. In conclusion, to preserve the clarity of the functions of each module and to enable efficient supervised learning, we have chosen $\alpha=1.0$ as the loss weight.
IV-B3 Denoising module architecture
Table III presents an ablation study that demonstrates the impact of the architecture of the denoising module. We evaluate different denoising module configurations with diverse scales and blocks to determine their efficiency. According to Table III, the variant that employs a multi-scale architecture consistently outperforms the single-scale architecture in terms of PSNR and SSIM metrics. In addition, increasing the number of blocks also improves denoising performance. By evaluating factors such as parameters and indices, our HDI-PRNet is designed with three scales and two blocks to implement the denoising module. Compared with the baseline model with a single scale and a single block, we see a PSNR improvement of 0.1 dB.
IV-B4 Deblurring module architecture
We continue our exploration by analyzing how different orders of the Neumann series expansion, specifically at orders $1$, $3$, and $5$, influence the reconstruction performance. Table IV illustrates an improvement in performance with more expansion terms, which confirms the effectiveness of the expansion architecture. Therefore, we choose an expansion of order $5$ in the deblurring module for the optimal configuration.
IV-B5 DDLB architecture
We evaluate the performance of DDLB by carrying out comparative analyses with several alternative variants. Table V displays the PSNR and SSIM metrics with a super-resolution scaling factor of ×$\times 2$. It is observed that focusing solely on either the spatial or frequency branch presents difficulties in efficiently recovering the image from blur degradation. The DEA module integrates these two branches more effectively than a straightforward addition. Compared to the baseline module in the spatial branch, the PSNR achieves an improvement of 0.2dB.
TABLE VI: Comparison of different SRx2 models on various scene classes of WHU-RS19 (PSNR ($dB$)/SSIM). The best and second-best results are highlighted in Bold font and underlined ones, respectively.
WHU-RS19 Airport Beach Bridge Commercial Desert Farmland FootballField Forest Industrial Meadow
Methods PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
CTNet [48] 27.08/0.737 34.41/0.944 31.79/0.854 24.31/0.693 36.79/0.898 33.52/0.824 28.05/0.793 27.03/0.586 25.54/0.718 34.81/0.815
HSENet [49] 27.55/0.745 34.41/0.950 32.25/0.858 24.60/0.694 36.83/0.901 33.95/0.828 28.39/0.799 27.18/0.594 25.94/0.723 34.84/0.813
D-DBPN [45] 27.80/0.758 35.67/0.945 32.54/0.863 24.74/0.704 37.03/0.900 33.82/0.828 28.76/0.809 27.39/0.607 26.12/0.734 34.95/0.815
RDN [17] 27.90/0.761 36.18/0.948 32.65/0.867 24.79/0.706 37.29/0.903 34.16/0.835 28.89/0.818 27.44/0.612 26.22/0.739 35.06/0.817
TransENet [28] 27.92/0.762 36.28/0.946 32.54/0.867 24.84/0.708 37.33/0.903 33.99/0.832 29.09/0.820 27.47/0.617 26.37/0.745 34.96/0.814
RCAN [10] 27.97/0.765 36.73/0.949 32.83/0.870 24.94/0.714 37.43/0.904 34.28/0.838 29.14/0.823 27.41/0.612 26.41/0.747 35.15/0.820
SAN [20] 27.98/0.761 36.64/0.943 32.49/0.861 24.87/0.709 37.22/0.903 34.27/0.835 28.97/0.817 27.45/0.612 26.29/0.740 35.12/0.819
MHAN [24] 27.91/0.762 36.49/0.951 32.58/0.865 24.88/0.711 37.38/0.904 34.23/0.835 28.96/0.820 27.43/0.613 26.37/0.743 35.02/0.816
SwinIR [12] 27.96/0.764 37.18/0.950 32.62/0.868 24.88/0.712 37.44/0.904 34.24/0.836 29.00/0.823 27.36/0.612 26.33/0.747 35.13/0.819
HAUNet [27] 28.05/0.762 36.60/0.948 32.81/0.870 24.97/0.711 37.50/0.904 34.40/0.838 29.24/0.823 27.50/0.619 26.53/0.747 35.17/0.820
TTST [26] 27.97/0.763 36.59/0.944 32.72/0.869 24.93/0.712 37.52/0.905 34.26/0.835 29.05/0.821 27.39/0.611 26.38/0.746 35.11/0.819
CAT [46] 28.10/0.768 37.15/0.954 32.93/0.871 25.11/0.720 37.53/0.905 34.36/0.838 29.42/0.832 27.57/0.626 26.56/0.755 35.15/0.819
RGT [47] 28.19/0.771 37.18/0.949 32.83/0.872 25.09/0.719 37.63/0.906 34.50/0.841 29.37/0.831 27.50/0.624 26.56/0.754 35.21/0.821
DAT [22] 28.25/0.772 36.99/0.948 32.83/0.873 25.11/0.721 37.65/0.906 34.53/0.841 29.48/0.832 27.52/0.624 26.60/0.756 35.20/0.821
HDI-PRNet 28.55/0.780 37.80/0.950 33.48/0.878 25.44/0.731 37.58/0.906 34.59/0.845 29.73/0.836 27.84/0.634 27.08/0.767 35.22/0.822
WHU-RS19 Mountain Park Parking Pond Port RailwayStation Residential River Viaduct Average
Methods PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
CTNet [48] 24.46/0.579 27.44/0.701 25.81/0.805 35.10/0.902 26.62/0.806 24.56/0.621 24.20/0.724 27.63/0.699 25.17/0.693 28.64/0.757
HSENet [49] 24.66/0.582 27.84/0.704 26.52/0.817 35.57/0.905 27.03/0.811 24.81/0.623 24.48/0.730 27.77/0.703 25.45/0.701 28.94/0.762
D-DBPN [45] 24.81/0.592 27.91/0.708 26.92/0.830 35.56/0.901 27.31/0.821 25.12/0.644 24.74/0.742 27.94/0.710 25.96/0.725 29.20/0.770
RDN [17] 24.83/0.594 27.99/0.711 26.93/0.833 35.82/0.907 27.31/0.824 25.19/0.653 24.74/0.745 28.01/0.714 26.04/0.732 29.33/0.774
TransENet [28] 24.87/0.604 28.03/0.713 26.94/0.836 35.65/0.907 27.33/0.825 25.08/0.646 24.82/0.748 27.98/0.719 26.09/0.733 29.33/0.776
RCAN [10] 24.92/0.602 28.09/0.716 27.01/0.838 35.97/0.909 27.39/0.827 25.21/0.658 24.90/0.752 28.02/0.719 26.24/0.740 29.46/0.779
SAN [20] 24.87/0.599 28.05/0.712 26.92/0.833 35.82/0.907 27.35/0.825 25.17/0.650 24.83/0.747 28.09/0.716 26.03/0.729 29.38/0.774
MHAN [24] 24.89/0.600 28.07/0.713 27.14/0.837 35.81/0.906 27.48/0.828 25.19/0.650 24.89/0.750 28.10/0.718 26.13/0.733 29.40/0.776
SwinIR [12] 24.88/0.604 28.05/0.716 26.94/0.839 35.91/0.907 27.38/0.829 25.01/0.652 24.87/0.751 27.94/0.718 26.13/0.738 29.42/0.778
HAUNet [27] 25.00/0.607 28.10/0.715 27.31/0.844 36.03/0.910 27.43/0.828 25.29/0.656 24.90/0.751 28.09/0.720 26.25/0.736 29.52/0.779
TTST [26] 24.86/0.600 28.07/0.714 27.13/0.840 35.91/0.909 27.43/0.829 25.31/0.651 24.92/0.752 28.07/0.719 26.09/0.733 29.44/0.777
CAT [46] 24.95/0.608 28.18/0.721 27.46/0.850 35.99/0.909 27.70/0.834 25.38/0.661 25.10/0.761 28.18/0.725 26.46/0.746 29.63/0.784
RGT [47] 24.95/0.610 28.25/0.722 27.27/0.845 36.02/0.910 27.67/0.834 25.31/0.662 25.10/0.761 28.15/0.725 26.46/0.749 29.63/0.784
DAT [22] 25.08/0.615 28.25/0.722 27.39/0.847 36.16/0.911 27.76/0.836 25.50/0.668 25.16/0.763 28.22/0.728 26.53/0.752 29.68/0.785
HDI-PRNet 25.27/0.627 28.48/0.727 28.18/0.860 36.24/0.911 28.35/0.845 26.05/0.693 25.58/0.772 28.54/0.735 26.87/0.761 30.03/0.793
TABLE VII: Comparison of different SRx3 models on various scene classes of WHU-RS19 (PSNR ($dB$)/SSIM). The best and second-best results are highlighted in Bold font and underlined ones, respectively.
WHU-RS19 Airport Beach Bridge Commercial Desert Farmland FootballField Forest Industrial Meadow
Methods PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
CTNet [48] 26.10/0.716 34.47/0.945 29.89/0.843 23.08/0.636 36.60/0.895 32.52/0.800 26.78/0.757 26.64/0.556 24.47/0.665 34.10/0.802
HSENet [49] 26.59/0.726 35.66/0.942 30.47/0.851 23.36/0.645 36.78/0.896 33.00/0.808 27.22/0.769 26.75/0.562 24.89/0.679 34.18/0.803
D-DBPN [45] 26.82/0.739 35.57/0.948 31.01/0.860 23.96/0.673 36.76/0.894 32.92/0.809 27.77/0.785 26.95/0.580 25.44/0.702 34.40/0.807
RDN [17] 26.90/0.741 36.40/0.951 31.07/0.861 23.89/0.671 37.01/0.899 33.26/0.813 27.73/0.790 26.93/0.578 25.39/0.701 34.42/0.807
TransENet [28] 27.09/0.744 36.91/0.953 31.06/0.864 23.94/0.674 36.86/0.899 33.18/0.813 27.79/0.790 27.01/0.587 25.49/0.704 34.34/0.806
RCAN [10] 27.09/0.748 36.57/0.948 31.18/0.865 23.99/0.680 37.03/0.899 33.29/0.814 27.87/0.794 27.07/0.593 25.49/0.707 34.53/0.810
SAN [20] 26.86/0.741 35.83/0.944 30.90/0.861 23.88/0.671 36.95/0.898 33.20/0.812 27.68/0.785 26.91/0.578 25.41/0.701 34.44/0.807
MHAN [24] 27.01/0.745 36.20/0.951 31.18/0.864 23.99/0.676 36.99/0.898 33.26/0.814 27.84/0.791 26.99/0.585 25.53/0.706 34.46/0.809
SwinIR [12] 27.10/0.749 37.11/0.954 31.21/0.866 24.00/0.680 37.08/0.900 33.30/0.814 27.81/0.795 27.06/0.592 25.53/0.709 34.48/0.809
HAUNet [27] 27.09/0.745 36.61/0.955 31.21/0.864 23.98/0.673 37.02/0.899 33.38/0.815 27.88/0.788 27.09/0.588 25.56/0.704 34.46/0.807
TTST [26] 27.00/0.743 35.63/0.955 31.13/0.862 23.92/0.671 36.69/0.898 33.31/0.814 27.85/0.787 27.00/0.582 25.47/0.701 34.46/0.808
CAT [46] 27.35/0.755 36.94/0.949 31.37/0.868 24.16/0.687 37.13/0.900 33.45/0.818 28.16/0.802 27.13/0.595 25.69/0.715 34.58/0.810
RGT [47] 27.31/0.755 37.29/0.955 31.37/0.869 24.16/0.687 37.14/0.900 33.47/0.819 28.19/0.804 27.13/0.596 25.69/0.716 34.54/0.809
DAT [22] 27.34/0.756 37.07/0.953 31.53/0.870 24.20/0.689 37.15/0.900 33.51/0.820 28.26/0.804 27.14/0.598 25.75/0.718 34.56/0.810
HDI-PRNet 27.53/0.759 37.62/0.957 32.02/0.874 24.41/0.692 36.97/0.900 33.66/0.822 28.61/0.808 27.23/0.596 26.02/0.723 34.60/0.811
WHU-RS19 Mountain Park Parking Pond Port RailwayStation Residential River Viaduct Average
Methods PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
CTNet [48] 23.53/0.501 26.72/0.653 24.04/0.735 32.81/0.883 25.21/0.785 23.87/0.575 22.79/0.660 26.57/0.648 24.22/0.638 27.59/0.720
HSENet [49] 23.74/0.514 27.05/0.661 24.59/0.749 33.12/0.888 25.48/0.792 24.31/0.596 23.17/0.673 26.93/0.658 24.59/0.653 27.98/0.729
D-DBPN [45] 23.93/0.525 27.31/0.672 25.27/0.775 33.36/0.890 26.06/0.805 24.78/0.623 23.57/0.693 27.14/0.668 25.10/0.682 28.31/0.743
RDN [17] 23.95/0.533 27.35/0.673 25.19/0.776 33.46/0.893 25.97/0.806 24.80/0.626 23.46/0.692 27.19/0.669 25.08/0.684 28.38/0.745
TransENet [28] 23.97/0.540 27.38/0.675 25.54/0.788 33.32/0.893 25.98/0.808 24.91/0.628 23.63/0.700 27.23/0.675 25.09/0.686 28.44/0.748
RCAN [10] 24.03/0.544 27.46/0.679 25.17/0.784 33.55/0.895 26.03/0.812 24.96/0.640 23.51/0.699 27.30/0.678 25.28/0.696 28.48/0.751
SAN [20] 23.91/0.530 27.35/0.673 25.12/0.774 33.41/0.892 25.89/0.805 24.73/0.623 23.50/0.694 27.14/0.668 25.03/0.679 28.31/0.743
MHAN [24] 24.00/0.538 27.43/0.676 25.23/0.782 33.57/0.894 26.09/0.810 24.94/0.634 23.58/0.698 27.25/0.673 25.21/0.688 28.45/0.748
SwinIR [12] 24.05/0.547 27.44/0.679 25.44/0.790 33.55/0.895 26.14/0.814 24.97/0.637 23.61/0.703 27.29/0.677 25.26/0.696 28.53/0.752
HAUNet [27] 24.04/0.539 27.35/0.671 25.64/0.792 33.54/0.893 26.07/0.808 24.93/0.627 23.65/0.700 27.33/0.675 25.16/0.685 28.51/0.748
TTST [26] 23.98/0.535 27.40/0.674 25.39/0.779 33.41/0.892 26.04/0.807 24.77/0.621 23.60/0.698 27.26/0.674 25.17/0.686 28.38/0.746
CAT [46] 24.13/0.553 27.47/0.681 25.80/0.804 33.59/0.896 26.33/0.818 25.18/0.647 23.77/0.710 27.38/0.683 25.47/0.707 28.67/0.757
RGT [47] 24.09/0.550 27.57/0.683 25.66/0.799 33.73/0.897 26.34/0.819 25.25/0.652 23.77/0.710 27.41/0.684 25.47/0.708 28.70/0.758
DAT [22] 24.12/0.552 27.53/0.682 25.89/0.803 33.71/0.897 26.37/0.819 25.18/0.651 23.79/0.712 27.44/0.684 25.52/0.709 28.72/0.759
HDI-PRNet 24.24/0.550 27.65/0.684 26.59/0.818 33.74/0.896 26.82/0.825 25.55/0.662 24.31/0.722 27.53/0.685 25.66/0.711 28.96/0.762
TABLE VIII: Comparison of different SRx4 models on various scene classes of WHU-RS19 (PSNR ($dB$)/SSIM). The best and second-best results are highlighted in Bold font and underlined ones, respectively.
WHU-RS19 Airport Beach Bridge Commercial Desert Farmland FootballField Forest Industrial Meadow
Methods PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
CTNet [48] 24.88/0.655 35.59/0.938 30.01/0.835 22.09/0.552 35.71/0.892 32.09/0.790 25.39/0.701 25.56/0.470 23.62/0.605 33.58/0.797
HSENet [49] 25.64/0.695 37.12/0.947 30.86/0.851 22.67/0.598 36.25/0.897 32.75/0.804 26.28/0.738 26.02/0.511 24.48/0.657 33.92/0.800
D-DBPN [45] 25.59/0.694 36.30/0.942 30.92/0.851 22.72/0.599 36.06/0.895 32.62/0.802 26.24/0.737 25.97/0.506 24.46/0.655 33.86/0.800
RDN [17] 25.52/0.693 36.85/0.945 30.85/0.852 22.62/0.595 36.21/0.897 32.72/0.803 26.15/0.738 25.93/0.509 24.36/0.652 33.89/0.800
TransENet [28] 25.69/0.696 37.40/0.948 30.88/0.854 22.73/0.600 36.03/0.897 32.74/0.804 26.39/0.745 26.08/0.517 24.54/0.659 33.92/0.801
RCAN [10] 25.68/0.702 37.07/0.946 30.93/0.855 22.76/0.606 36.20/0.897 32.86/0.806 26.35/0.746 26.06/0.516 24.61/0.665 33.90/0.800
SAN [20] 25.53/0.692 36.65/0.944 30.81/0.850 22.61/0.593 36.19/0.896 32.69/0.802 26.11/0.733 25.94/0.506 24.40/0.651 33.90/0.799
MHAN [24] 25.64/0.696 37.20/0.947 30.69/0.851 22.70/0.598 36.24/0.897 32.75/0.803 26.27/0.740 25.98/0.512 24.49/0.656 33.93/0.800
SwinIR [12] 25.79/0.704 37.49/0.949 30.99/0.855 22.81/0.609 36.31/0.898 32.90/0.807 26.47/0.750 26.12/0.521 24.66/0.668 33.96/0.801
HAUNet [27] 25.67/0.695 37.16/0.947 30.87/0.852 22.71/0.598 36.17/0.897 32.89/0.806 26.42/0.742 26.11/0.518 24.55/0.658 33.95/0.800
TTST [26] 25.76/0.703 37.21/0.948 30.85/0.854 22.80/0.607 36.23/0.898 32.89/0.806 26.46/0.748 26.09/0.519 24.68/0.667 33.93/0.801
CAT [46] 25.84/0.706 37.15/0.947 31.01/0.855 22.85/0.611 36.33/0.898 33.00/0.809 26.62/0.755 26.11/0.521 24.78/0.673 34.04/0.802
RGT [47] 25.93/0.710 37.60/0.949 31.25/0.858 22.91/0.613 36.37/0.898 33.05/0.810 26.72/0.758 26.15/0.523 24.85/0.675 34.06/0.802
DAT [22] 25.92/0.710 37.54/0.949 31.28/0.858 22.92/0.614 36.41/0.898 33.05/0.810 26.71/0.758 26.16/0.522 24.87/0.676 34.06/0.802
HDI-PRNet 26.10/0.712 36.88/0.948 31.61/0.862 23.05/0.615 36.40/0.898 33.09/0.811 26.87/0.757 26.19/0.526 24.94/0.674 34.07/0.803
WHU-RS19 Mountain Park Parking Pond Port RailwayStation Residential River Viaduct Average
Methods PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
CTNet [48] 23.01/0.433 25.38/0.585 23.10/0.676 30.22/0.838 24.27/0.736 23.05/0.501 21.79/0.586 25.98/0.606 23.08/0.562 26.74/0.671
HSENet [49] 23.43/0.482 25.87/0.612 24.30/0.738 30.81/0.850 25.04/0.771 23.96/0.570 22.62/0.643 26.54/0.641 24.13/0.631 27.49/0.707
D-DBPN [45] 23.37/0.474 25.81/0.612 24.29/0.738 30.72/0.848 25.14/0.771 23.92/0.569 22.64/0.640 26.48/0.637 24.20/0.635 27.42/0.705
RDN [17] 23.39/0.479 25.83/0.612 24.16/0.733 30.74/0.848 25.01/0.769 23.96/0.572 22.52/0.637 26.47/0.638 24.16/0.635 27.42/0.705
TransENet [28] 23.48/0.488 25.92/0.615 24.52/0.752 30.83/0.851 25.20/0.775 24.05/0.572 22.77/0.651 26.57/0.644 24.23/0.638 27.56/0.710
RCAN [10] 23.46/0.489 25.94/0.617 24.39/0.752 30.82/0.852 25.14/0.777 24.10/0.583 22.68/0.652 26.55/0.644 24.37/0.647 27.55/0.713
SAN [20] 23.37/0.474 25.81/0.611 24.11/0.731 30.72/0.848 24.96/0.768 23.87/0.564 22.53/0.637 26.42/0.636 24.03/0.623 27.38/0.702
MHAN [24] 23.42/0.481 25.89/0.614 24.19/0.740 30.78/0.850 25.08/0.772 23.96/0.571 22.59/0.644 26.50/0.640 24.18/0.634 27.48/0.707
SwinIR [12] 23.50/0.492 25.95/0.619 24.75/0.762 30.87/0.853 25.29/0.780 24.15/0.583 22.75/0.657 26.62/0.648 24.36/0.648 27.65/0.715
HAUNet [27] 23.46/0.482 25.87/0.612 24.71/0.760 30.80/0.851 25.17/0.774 24.04/0.569 22.74/0.652 26.59/0.645 24.22/0.633 27.56/0.709
TTST [26] 23.50/0.491 25.94/0.617 24.67/0.756 30.90/0.853 25.28/0.779 24.09/0.580 22.73/0.654 26.61/0.647 24.34/0.646 27.61/0.714
CAT [46] 23.50/0.489 25.96/0.620 24.91/0.770 30.91/0.854 25.38/0.782 24.27/0.589 22.86/0.662 26.64/0.649 24.48/0.656 27.70/0.718
RGT [47] 23.54/0.496 26.06/0.622 25.01/0.772 30.99/0.854 25.45/0.785 24.33/0.594 22.94/0.664 26.70/0.651 24.56/0.658 27.79/0.720
DAT [22] 23.55/0.495 26.05/0.622 25.06/0.772 30.97/0.855 25.43/0.784 24.34/0.594 22.92/0.664 26.70/0.653 24.54/0.659 27.79/0.720
HDI-PRNet 23.62/0.498 26.21/0.623 25.29/0.778 31.09/0.854 25.58/0.783 24.52/0.601 23.21/0.668 26.74/0.652 24.60/0.657 27.87/0.721
Figure 4: PSNR violin-plot of different x2 SR methods on WHU-RS19 dataset.
IV-C Comparisons with other methods on WHU-RS19 dataset
In this section, we assess the effectiveness of our proposed method by comparing it with popular image restoration methods using the WHU-RS19 dataset. Tables VI, VII, and VIII present the quantitative analysis of various evaluation metrics on different super-resolution scales. Our proposed HDI-PRNet achieves the best reconstruction performance in terms of average reconstruction metrics, showing average improvements of 0.35dB, 0.24dB, and 0.08dB. Furthermore, it also produces optimal or suboptimal reconstruction results in most land-use categories.
Fig. 3 shows a visualization comparison of the restoration results obtained by different methods in the WHU-RS19 dataset. Our method significantly alleviates the degradation caused by noise, blur, and downsampling, resulting in sharper images that are visually close to the ground truth. Although other methods have attained some success, they cannot fully restore the details of texture and structure. Specifically, the restoration results of CTNet and HSENet have many artifacts. Approaches such as DAT and HAUNet have complex designed architectures that can extract richer features, but they all produce over-smoothing results. In contrast, our method not only removes noise but also effectively restores image details and edges, obtaining the clearest and most accurate remote sensing images.
Fig. 4 displays a violin-plot that serves as a tool for assessing the stability in performance of the high-order restoration framework. The plot includes three primary lines: the first quartile (Q1), the median, and the third quartile (Q3), which are useful for assessing the result’s dispersion. The Q1, median and Q3 values of the proposed method are the highest, indicating better overall performance compared to other approaches. The curve in the violin-plot represents a kernel density map, reflecting the data distribution shape. As depicted in Fig. 4, our method exhibits a more concentrated data distribution with less variance, contributing to greater model stability than alternative methods.
Figure 5: Visualization comparison of different methods on the real satellite images.
TABLE IX: Comparison of different restoration models on real satellite images.
BRISQUE HSENet TransENet MHAN MHAN (Bicubic) SwinIR SwinIR (Bicubic) TTST HAUNet DAT HDI-PRNet
Jilin-1 41.3774 41.1180 35.8704 43.1953 37.1765 43.6537 37.3207 43.4201 36.8362 36.5810
Gaofen-1 40.3188 39.6849 40.2469 45.5903 41.4091 45.4889 40.9722 41.3830 40.7827 39.6351
Gaofen-2 37.3217 38.3410 37.9373 39.8892 39.3446 39.4106 39.2308 39.1100 38.5144 35.9592
Gaofen-6 36.8403 38.3920 38.3865 41.4711 38.9494 41.1936 38.7809 38.9542 38.4101 36.7054
Gaofen-7 37.7147 38.0881 37.7705 43.6968 38.6543 43.0609 38.3997 37.8802 37.1173 37.0412
IV-D Comparisons on real remote sensing images
To validate the generalizability of the proposed method, we perform experiments using real remote sensing images captured by the Jilin-1, Gaofen-1, Gaofen-2, Gaofen-6, and Gaofen-7 satellites. To facilitate testing, the dataset includes $824$ Jilin-1 images that have been resized to $64\times 64$ pixels, along with $400$ images from each Gaofen satellite resized to $128\times 128$ pixels.
The visualization comparison is shown in Fig.5. MHAN, SwinIR, DAT, and TTST have evidently enhanced image quality to a certain extent; however, they have excessively smoothed the original image, unintentionally removing some details by mistaking them for noise or artifacts. In comparison to other methods, the reconstructed image using the proposed method is significantly clearer, particularly in detail and texture, restoring these aspects more accurately and closely to the original image without producing artifacts or over-smoothing. Furthermore, we train the MHAN and SwinIR models employing bicubic degradation and evaluate them on real remote sensing datasets. Both MHAN (bicubic) and SwinIR (bicubic) demonstrate a slight but not evident improvement in the quality of real remote sensing images, which further demonstrates the rationality of our model’s architecture and configurations.
Since real remote sensing images lack HR ground truth, we employ BRISQUE to evaluate the predicted results. As shown in Table IX, HDI-PRNet achieves the best or second-best results on real images from five different satellites. Although MHAN records the highest score on Jilin-1, its predictions are visually over-smoothed, which is not preferable. Overall, these significant improvements in both visual quality and reconstruction performance metrics strongly show the effectiveness and practical utility of HDI-PRNet to restore real remote sensing images, further highlighting the excellence of our proposed HDI-PRNet.
V Conclusion
In this paper, we proposed a novel remote sensing high-order restoration framework. Our framework is based on the inverse design of the high-order degradation model to gradually restore the image, which provides a theoretical basis for the design of the DL framework. In addition, we introduced a denoising module that simulates the proximal mapping operator to learn the denoising prior representation of the current image. To solve the deblurring problem, we proposed a deblurring module that ensures the accurate representation of the degradation operator based on the expansion of the Neumann series and dual-domain degradation learning. We also introduced an intermediate loss in the middle of the network so that each module has supervision and the overall network is more interpretable. We performed several high-order degradation reconstruction experiments on WHU-RS19 and real remote sensing images. Experimental results show that our HDI-PRNet achieves competitive results on both synthetic and real images.
In the future, we believe that our method can be extended to other image restoration problems to perform more complex and diverse image restoration tasks, such as image dehazing and image artifact removal, and obtain better restoration performance on real images.